Suppose the demand for a good is modeled by the equation Q = 5 - 3P, and the supply for the good is modeled by Q = 2 + P. Fill in the blanks below with decimal answers (i.e. no fractions).

The equilibrium quantity of the good will be ________ and the equilibrium price will be ________.
Answer #1

Setting QD=QS to find the equilibrium quantity and price

5-3P=2+P

5-2=4P

2=4P

Equilibrium Price=2/4 =0.5

Equilibrium Quantity=5-3(0.5) = 3.5